# SN 2021foa: Deriving a continuity between SN IIn and SN Ibn

### v1

@@ 2.4 Distance & extinction
Adopting H0 = 73 km s-1 Mpc-1, m = 0.27 and Λ = 0.73, we obtain a distance of 34.89 2.44 Mpc () corrected for Virgo, Shapley and GA (corresponding to a redshift 0.00839 555https://ned.ipac.caltech.edu/byname?objname=IC0863&hconst=73&omegam=0.27&omegav=0.73&wmap=1&corr_z=2) for SN 2021foa. This value is the same as that adopted by R22. The Milky Way extinction along the line of sight of SN 2021foa is mag (Schlafly & Finkbeiner, 2011). We see a conspicuous dip at 5892.5 Å from Na 1D in the spectra of SN 2021foa taken on 2021-03-20, 2021-03-23 and 2021-03-25. For estimating the extinction due to the host galaxy, we estimate equivalent widths of the Na 1D line iteratively three times in the combined spectra of these three dates to increase the signal-to-noise ratio. Using the formulation by Poznanski et al. (2012), we estimate host galaxy = 0.40 0.19 mag. We multiply this reddening value by 0.86 to be consistent with the recalibration of Milky Way extinction by Schlafly & Finkbeiner (2011). Thus, we adopt a total extinction of = 0.57 0.16 mag. We use these values of distance and extinction throughout the paper, which is also consistent with the values quoted by R22.
@@ 3 Spectroscopic Evolution
We conducted the spectroscopic follow-up of SN 2021foa from 10.8 d to 58.6 d post -band maximum. The complete spectral evolution of SN 2021foa, which includes our data and those published in R22, are shown in Figure 1. The early time spectral sequence shows prominent lines of H (6563 Å) along with H (4861 Å) and H (4340 Å). From 10.8 d to d, the He (5876 Å) line also starts developing but is not very prominent. The H profile initially shows narrow emission, but develops a narrow P-Cygni feature at around 5.8 d in our spectra. The spectra from R22 show an even earlier appearance of this narrow P-Cygni feature, which can be attributed to the higher resolution of their spectrum. The H also develops this narrow P-Cygni feature in the 3.2 d spectra. The early spectral sequence till d does not show prominent lines of Fe ii (4924, 5018, and 5169 Å), which are characteristic of a typical SN ejecta.
Post d, the He i (5876 Å) line becomes prominent. The other He i (6678, 7065 Å) lines start developing at this phase. The narrow P-Cygni on top of H is clearly seen with the blue wing extending up to 1900 km s-1. From 7.2 d to 17.2 d, the H and H show very complex profiles. This is also the phase where we see a second shoulder appearing in the light curve of SN 2021foa (c.f.r Section 4).
@@
From 26.8 d to 58.6 d, we see that both H and He i grow in strength. This also marks the phase where we see ejecta signatures in the spectral evolution. The [Ca ii] 7291, 7324 Å lines emerge at 7300 Å, which could also be blended with He i 7281 Å. This phase also marks the appearance of the broad Ca II NIR triplet. At late times ( d), the intermediate to broad-width H-lines develop a slight blueshift in the observed profiles.
Overall, the spectral behaviour shows a striking similarity with an interacting SNe IIn early on, which later on is overtaken by a SN Ibn like behaviour with He lines. The very late spectra shows prominent ejecta signatures of Ca.
@@ 3.1 Line luminosities and line ratios:
To study whether SN 2021foa belongs to the SNe IIn or the SNe Ibn regime, we compare the evolution of H and He i line luminosities with the sample. The line luminosities are estimated by integrating the continuum-subtracted line regions of the de-reddened spectra. The top panel of the Figure 2 shows the H luminosities of SNe IIn (blue) and SNe Ibn (pink) sample and SN 2021foa marked in black. The H luminosities of SNe IIn and SNe Ibn are well separated in luminosity scales. SN 2021foa shows similarity with SN 2011hw in the H space, which is an SN Ibn with a significant amount of residual Hydrogen (Pastorello et al., 2015a). On the contrary, in the He i luminosity scale (middle panel of Figure 2), SNe IIn and SNe Ibn do not show a clear distinction. The He i 5876 Å luminosity of SN 2021foa matches with SN 2006jc (Pastorello et al., 2007) over the evolution. The lower panel of Figure 2 shows the luminosity ratios of He i 5876 Å to H. The distinction between these two classes of objects is the most prominent in this plot, however, we want to mention that there might be some residual H luminosity from the host galaxy in case of SNe Ibn.
@@
Figure 3 aims to show how much the H and He i luminosity contributes to the total optical luminosity (-band; Lv) at peak for a group of SNe IIn and SNe Ibn. To avoid possible biases in the distribution, we added more diverse SNe IIn and SNe Ibn sample for this comparison plot to highlight the position of SN 2021foa in the phase space. The additional SNe IIn used for this comparison plot are : SNe 1998S (Fassia et al., 2001), 2005ip (Stritzinger et al., 2012), 2006gy (Smith & McCray, 2007; Agnoletto et al., 2009), 2006tf (Smith et al., 2008b), 2007od (Andrews et al., 2010), 2009kn (Kankare et al., 2012), 2011ht (Mauerhan et al., 2013), PTF11oxu/2011jc (Nyholm et al., 2020), ASASSN-14il (Dukiya et al., 2024), 2015da (Smith et al., 2024) and ASASSN-15ua (Dickinson et al., 2024). The additional SNe Ibn used for the comparison are : SNe OGLE-2012-SN-006 (Pastorello et al., 2015b), LSQ12btw (Pastorello et al., 2015c), LSQ13ccw (Pastorello et al., 2015c), ASASSN-15ed (Pastorello et al., 2015d), SNe 2014av (Pastorello et al., 2016), 2015U (Shivvers et al., 2016) and 2015G (Hosseinzadeh et al., 2017). We added SNe IIn/Ibn of different CSM configurations, long-lived, short-lived, with and without precursor to diversify this plot. Also, we chose only those SNe for which V-band observations around maximum are available with a good signal-to-noise ratio spectrum of min 10. We plot the ratio of H to the peak -band optical luminosity against the ratio of He i to the peak -band optical luminosity for SN 2021foa and the comparison sample. However, these SNe are in general have asymmetric CSM geometries (Smith, 2017; Fraser, 2020) and the peak -band line luminosities may be slightly affected by our viewing angle. We see that the SNe IIn and SNe Ibn in our sample are well separated in this phase space, with SN 2021foa (black star) lying in between the two sub-classes around peak luminosity. SN 2021foa shares remarkable similarities with SN 2011hw in this space as well. Pastorello et al. (2015a) have shown that SN 2011hw is also a SN Ibn with significant residual H. The distributions of SNe IIn and SNe Ibn also help in deciphering the fact that it is the H contributing to the optical luminosity that demarcates the two classes. He i distribution is blended for the SNe IIn and SNe Ibn population. However, we want to remark that a statistically decent sample would help in further verification of this. Nonetheless, SN 2021foa clearly lies at the junction between the two populations at this phase.
@@ 3.2 Spectral Comparison
Figure 4 shows the pre-maximum spectral comparison of SN 2021foa with a group of SNe IIn and SNe Ibn. The pre-maximum spectral profile of SN 2021foa looks very similar to all the SNe IIn in our comparison sample. However, the H-lines are more prominent in the SNe IIn compared to SN 2021foa. In contrast, the SNe Ibn show little to no hydrogen in their spectra. In SN 2011hw, an SN Ibn with significant residual hydrogen, both H and He lines are visible at this phase, while for SN 2021foa He i lines are not seen. Most SNe Ibn at this phase show flash features, which is absent in our observed profile. Overall, at this phase, SN 2021foa behaves more like a SN IIn.
Figure 5 shows the spectral comparison at 20 d after the -maximum. This phase marks the remarkable transition of SN 2021foa, showing prominent lines of H and He i simultaneously. We also see narrow P-Cygni of H in SN 2021foa at this phase, similar to SNe 1996al and 2016jbu. The strength of H remains lower in strength for SN 2021foa than other SNe IIn; however, He i shows similar strength with most SNe Ibn in the sample. This separates SN 2021foa with the SNe IIn and SNe Ibn population, again justifying our case of SN 2021foa having strong H and He i emission simultaneously at mid epochs and at similar strengths.
We also show the late-time spectral comparison of SN 2021foa in Figure 6. The spectral evolution of SN 2021foa at this phase matches very well with SNe 2005la and 2011hw, which are SNe Ibn with residual Hydrogen. SNe Ibn at this phase shows very strong He i lines, unlike SN 2021foa. Similarly, the He i lines are more prominent than traditional long-lasting SNe IIn. The H and He i line profiles of SN 2021foa also show similarities with SN 1996al at this phase, in addition to the similarities in the line-luminosity ratio, noted earlier.
@@ 4 Photometric Evolution
After the precursor, the lightcurve rose to peak in most of the optical bands in 6 d - 9 d which is consistent with SNe Ibn (Hosseinzadeh et al., 2017) and fast rising sample of SNe IIn (Nyholm et al., 2020). From the peak to about 14 d, it did not have much evolution and changed by only 0.3 - 0.5 mag in the optical wavebands. At 14 d, we see a shoulder in the light curve of SN 2021foa. This also marks the phase where we see the He i features developing and appearance of a narrow P-Cygni on top of H (c.f.r Section 3). The bump or the shoulder in the lightcurve of SN 2021foa is weaker in the redder bands than in the bluer bands. After this phase, the lightcurve drops sharply at a decline rate of about 3 mag in 50 days.
@@
Post 75 d, the lightcurve shows a flattening lasting from 75 d to 150 d post maximum. The flattening in the light curve of SN 2021foa was also noticed by R22. The late time flattening in the redder bands have been attributed to the formation of dust (Anupama et al., 2009), but, we do not have any NIR observations to verify this scenario (Smith et al., 2009; Benetti et al., 2016; Smith et al., 2024). The late time flattening could also be due to interaction with a uniform density CSM as we also see in our lightcurve modelling section (c.f.r Section 5.3). A recent paper by Farias (2024) shows existence of dust in the spectral evolution of SN 2021foa, justifying the evidence of newly formed dust/old dust in the ejecta-CSM front of SN 2021foa.
The double peak or hump seen in the lightcurve of SN 2021foa can also be reproduced overall by the grid of models by Suzuki et al. (2019) which are based on a CSM of mass 10 M⊙ assuming a disk-like geometry of the CSM.
@@
Figure 8 shows the (B-V)0 color evolution and the absolute magnitude lightcurve of SN 2021foa with other members of the SNe IIn and SNe Ibn subclass. The (B-V)0 color curve of SN 2021foa increases by 0.5 mag in colors from 10 d to about 40 d post maximum. In the phase space of color evolution, we see two sectors of events. One set of SNe (1996al, 2006jc, 2016jbu, 2018cnf; Category 1) in our plot, increases from red to blue from 10 d to about maximum and then becomes flatter in the color curve evolution while for the latter, we see a constant rising of red to blue colors upto 30-40 d post maximum (SNe; Category 2) and then drops in the color evolution. Our SN 2021foa follows the latter trend in the evolution. The SNe of Category 1 are the SNe IIn which had a long term precursor activity. Also, for these events the flattening in color evolution is seen after the second peak in the SN light curve (Brennan et al., 2022). This is in contrary to SN 2021foa which shows a short term precursor and becomes red upto 50 d post maximum. Post 50 d, the colorcurve again becomes blue and continues upto 150d. This marks also the phase where we see a change in the mass-loss rates of the evolution (see subsection 6).
@@
We compare the absolute magnitude (-band) lightcurve of SN 2021foa with a group of SNe IIn and SNe Ibn. For the cases where -band in not available, we use Johnson Cousin -band. The absolute magnitude lightcurve of SN 2021foa behaves similarly with other events having precursor activities. The precursor lightcurve (Event A) had an absolute magnitude 14 mag (R22) similar to SNe 2009ip and 2016jbu. The second peak in the lightcurve (Event B) lies fairly intermediate (MV = 17.8) among the SNe IIn and SNe Ibn (c.f.r Figure 1 of R22). Kiewe et al. (2012), through their studies have found have found that most SNe IIn with precursor events typically rises to second peak around 17 days. The event B typically rises to a maximum with absolute magnitude 18 mag 0.5 (Kiewe et al., 2012; Nyholm et al., 2020) followed by a bumpy decline. Our SN 2021foa also reaches a peak mag at around 20 days, however, SN 2021foa has a low luminosity compared to the sample of Kiewe et al. (2012); Nyholm et al. (2020). Overall, the light curve of SN 2021foa is similar in luminosity to both SNe IIn and SNe Ibn population with L 1042 - 1043 erg sec-1, but the lightcurve resembles more those of SNe IIn given the heterogeneity. The SNe Ibn in our comparison sample have instead more short-lived and less bumpy lightcurves, in accordance with the sample presented by Hosseinzadeh et al. (2017). Also, interestingly, some events with precursor activity like SNe 2009ip and 2018cnf show a light curve shoulder similar to that of SN 2021foa, 20 d post -band maximum. This is most likely associated with the change in the mass-loss rate happening years before explosion. This is also affected by the opacity effects influencing the lightcurve behaviour.

### v3

@@ 2.4 Distance & extinction
Adopting H0 = 73 km s-1 Mpc-1, = 0.27 and = 0.73, we obtain a distance of 34.89 2.44 Mpc ( = 32.71 0.15) corrected for Virgo, Shapley and GA (corresponding to a redshift z=0.00839 555https://ned.ipac.caltech.edu/byname?objname=IC0863&hconst=73&omegam=0.27&omegav=0.73&wmap=1&corr_z=2) for SN 2021foa. This value is the same as that adopted by R22. The Milky Way extinction along the line of sight of SN 2021foa is mag (Schlafly & Finkbeiner, 2011). We see a conspicuous dip at 5892.5 Å from Na 1D in the spectra of SN 2021foa taken on 2021-03-20, 2021-03-23 and 2021-03-25. For estimating the extinction due to the host galaxy, we estimate equivalent widths of the Na 1D line iteratively three times in the combined spectra of these three dates to increase the signal-to-noise ratio. Using the formulation by Poznanski et al. (2012), we estimate host galaxy = 0.40 0.19 mag. We multiply this reddening value by 0.86 to be consistent with the recalibration of Milky Way extinction by Schlafly & Finkbeiner (2011). Thus, we adopt a total extinction of = 0.57 0.16 mag. We use these values of distance and extinction throughout the paper, which is also consistent with the values quoted by R22.
@@ 3 Spectroscopic Evolution
We conducted the spectroscopic follow-up of SN 2021foa from 10.8 d to 58.6 d post -band maximum. The complete spectral evolution of SN 2021foa, which includes our data and those published in R22, are shown in Figure 1. The early time spectral sequence shows prominent lines of H 6563 Å along with H 4861 Å and H 4340 Å. From 10.8 d to d, the He 5876 Å line also starts developing but is not very prominent. The H profile initially shows narrow emission, but develops a narrow P-Cygni feature at around 5.8 d in our spectra. The spectra from R22 show an even earlier appearance of this narrow P-Cygni feature, which can be attributed to the higher resolution of their spectrum. The H also develops this narrow P-Cygni feature in the 3.2 d spectra. The early spectral sequence till d does not show prominent lines of Fe ii 4924, 5018, and 5169 Å, which are characteristic of a typical SN ejecta.
Post d, the He i 5876 Å line becomes prominent. The other He i 6678, 7065 Å lines start developing at this phase. The narrow P-Cygni on top of H is clearly seen with the blue wing extending up to 1900 km s-1. From 7.2 d to 17.2 d, the H and H show very complex profiles. This is also the phase where we see a second shoulder appearing in the light curve of SN 2021foa (c.f.r Section 4).
@@
From 26.8 d to 58.6 d, we see that both H and He i grow in strength. The narrow component of H seen at 44.2 d is due to poor resolution and is an artifact. This also marks the phase where we see ejecta signatures in the spectral evolution. The [Ca ii] 7291, 7324 Å lines emerge at 7300 Å, which could also be blended with He i 7281 Å. This phase also marks the appearance of the broad Ca II NIR triplet. At late times ( d), the intermediate to broad-width H-lines develop a slight blueshift in the observed profiles.
Overall, the spectral behavior shows a striking similarity with an interacting SNe IIn early on, which later on is overtaken by a SN Ibn like behaviour with He lines. The very late spectra shows ejecta signatures of Ca emerging along with H and He i.
@@ 3.1 Line luminosities and line ratios:
To study whether SN 2021foa belongs to the SNe IIn or the SNe Ibn regime, we compare the evolution of H and He i line luminosities with the sample. The line luminosities are estimated by integrating the continuum-subtracted line regions of the de-reddened spectra. The top panel of the Figure 2 shows the H luminosities of SNe IIn (blue) and SNe Ibn (pink) sample and SN 2021foa marked in black. The H luminosities of SNe IIn and SNe Ibn are well separated in luminosity scales. SN 2021foa shows similarity with SN 2011hw in the H space, which is an SN Ibn with a significant amount of residual Hydrogen (Pastorello et al., 2015a). On the contrary, in the He i luminosity scale (middle panel of Figure 2), SNe IIn and SNe Ibn do not show a clear distinction. The He i 5876 Å luminosity of SN 2021foa matches with SN 2006jc (Pastorello et al., 2007) over the evolution. The lower panel of Figure 2 shows the luminosity ratios of He i 5876 Å to H. The distinction between these two classes of objects is the most prominent in this plot, however, we want to mention that there might be some residual H luminosity from the host galaxy in case of SNe Ibn (possibly for SNe 2006jc, 2010al, 2014av and 2019uo from our sample; being located in the spiral arms of the galaxy.).
@@
Figure 3 aims to show how much the H and He i luminosity contributes to the total optical luminosity (-band; Lv) at peak for a group of SNe IIn and SNe Ibn. To avoid possible biases in the distribution, we added more diverse SNe IIn and SNe Ibn sample for this comparison plot to highlight the position of SN 2021foa in the phase space. The additional SNe IIn used for this comparison plot are : SNe 1998S (Fassia et al., 2001), 2005ip (Stritzinger et al., 2012), 2006gy (Smith & McCray, 2007; Agnoletto et al., 2009), 2006tf (Smith et al., 2008b), 2009kn (Kankare et al., 2012), 2011ht (Mauerhan et al., 2013), PTF11oxu/2011jc (Nyholm et al., 2020), ASASSN-14il (Dukiya et al., 2024), 2015da (Smith et al., 2024) and ASASSN-15ua (Dickinson et al., 2024). The additional SNe Ibn used for the comparison are : SNe OGLE-2012-SN-006 (Pastorello et al., 2015b), LSQ12btw (Pastorello et al., 2015c), LSQ13ccw (Pastorello et al., 2015c), ASASSN-15ed (Pastorello et al., 2015d), SNe 2014av (Pastorello et al., 2016), 2015U (Shivvers et al., 2016) and 2015G (Hosseinzadeh et al., 2017). We added SNe IIn/Ibn of different CSM configurations, long-lived, short-lived, with and without precursor to diversify this plot. Also, we chose only those SNe for which V-band observations around maximum are available with a good signal-to-noise ratio spectrum of min 10. We plot the ratio of H to the peak -band optical luminosity against the ratio of He i to the peak -band optical luminosity for SN 2021foa and the comparison sample. The surrounding axes shows the marginalized distribution with respect to the axes variable. However, these SNe are in general have asymmetric CSM geometries (Smith, 2017; Fraser, 2020) and the peak -band line luminosities may be slightly affected by our viewing angle. We see that the SNe IIn and SNe Ibn in our sample are well separated in this phase space, with SN 2021foa (black star) lying in between the two sub-classes around peak luminosity. SN 2021foa shares remarkable similarities with SN 2011hw in this space as well. Pastorello et al. (2015a) have shown that SN 2011hw is also a SN Ibn with significant residual H. The distributions of SNe IIn and SNe Ibn also help in deciphering the fact that it is the H contributing to the optical luminosity that demarcates the two classes. He i distribution is blended for the SNe IIn and SNe Ibn population. However, we want to remark that a statistically decent sample would help in further verification of this. Nonetheless, SN 2021foa clearly lies at the junction between the two populations at this phase.
@@ 3.2 Spectral Comparison
The first panel of Figure 4 shows the pre-maximum spectral comparison of SN 2021foa with a group of SNe IIn and SNe Ibn. The pre-maximum spectral profile of SN 2021foa looks very similar to all the SNe IIn in our comparison sample. However, the H-lines are more prominent in the SNe IIn compared to SN 2021foa. In contrast, the SNe Ibn show little to no hydrogen in their spectra. In SN 2011hw, an SN Ibn with significant residual hydrogen, both H and He lines are visible at this phase, while for SN 2021foa He i lines are not seen. Most SNe Ibn in our comparison sample at this phase show flash features, which is absent in our observed profile. Overall, at this phase, SN 2021foa behaves more like a SN IIn.
Middle panel of Figure 4 shows the spectral comparison at 20 d after the -maximum. This phase marks the remarkable transition of SN 2021foa, showing prominent lines of H and He i simultaneously. We also see narrow P-Cygni of H in SN 2021foa at this phase, similar to SNe 1996al and 2016jbu. The strength of H remains lower in strength for SN 2021foa than other SNe IIn; however, He i shows similar strength with most SNe Ibn in the sample. This separates SN 2021foa with the SNe IIn and SNe Ibn population, again justifying our case of SN 2021foa having strong H and He i emission simultaneously at mid epochs and at similar strengths.
We also show the late-time spectral comparison of SN 2021foa in the third panel of Figure 4. The spectral evolution of SN 2021foa at this phase matches very well with SNe 2005la and 2011hw, which are SNe Ibn with residual Hydrogen. SNe Ibn at this phase shows very strong He i lines, unlike SN 2021foa. Similarly, the He i lines are more prominent than traditional long-lasting SNe IIn. The H and He i line profiles of SN 2021foa also show similarities with SN 1996al at this phase, in addition to the similarities in the line-luminosity ratio, noted earlier.
@@ 4 Photometric Evolution
After the precursor, the lightcurve rose to peak in most of the optical bands in 6 d - 9 d which is consistent with SNe Ibn (Hosseinzadeh et al., 2017) and fast rising sample of SNe IIn (Nyholm et al., 2020). From the peak to about 14 d, it did not have much evolution and changed by only 0.3 - 0.5 mag in the optical wavebands. At 14 d, we see a shoulder in the light curve of SN 2021foa. This also marks the phase where we see the He i features developing and strengthening of a narrow P-Cygni on top of H (c.f.r Section 3). The bump or the shoulder in the lightcurve of SN 2021foa is weaker in the redder bands than in the bluer bands. After this phase, the lightcurve drops sharply at a decline rate of about 3 mag in 50 days.
@@
Post 75 d, the lightcurve shows a flattening lasting from 75 d to 150 d post maximum. The flattening in the light curve of SN 2021foa was also noticed by R22. The late time flattening in the redder bands have been attributed to the formation of dust (Anupama et al., 2009), but, we do not have any NIR observations to verify this scenario (Smith et al., 2009; Benetti et al., 2016; Smith et al., 2024). The late time flattening could also be due to interaction with a uniform density CSM as we also see in our lightcurve modelling section (c.f.r Section 5.3). The double peak or hump seen in the lightcurve of SN 2021foa can also be reproduced overall by the grid of models by Suzuki et al. (2019) which are based on a CSM of mass 10 M⊙ assuming a disk-like geometry of the CSM.
@@
Figure 6 shows the (B-V)0 color evolution and the absolute magnitude lightcurve of SN 2021foa with other members of the SNe IIn and SNe Ibn subclass. We have used near-simulatenous observations ( day) in B and V filter without any interpolations, to generate the color curve. The plotted epoch is the mean of observation epochs. The (B-V)0 color curve of SN 2021foa increases by 0.5 mag in colors from 10 d to about 40 d post maximum. In the phase space of color evolution, we see two sectors of events. One set of SNe (1996al, 2006jc, 2016jbu, 2018cnf; open symbols; Category 1) in our plot, increases from red to blue from 10 d to about maximum and then becomes flatter in the color curve evolution while for the latter, we see a constant rising of red to blue colors upto 30-40 d post maximum (SNe; closed symbols; Category 2) and then drops in the color evolution. For SNe 2019uo, 2019wep we cannot conclusively say anything due to limited data sets. Our SN 2021foa follows Category 2 in the evolution. The SNe of Category 1 are the SNe IIn which had a long term precursor activity. Also, for these events the flattening in color evolution is seen after the second peak in the SN light curve (Brennan et al., 2022). This is in contrary to SN 2021foa which shows a short term precursor and becomes red upto 50 d post maximum. Post 50 d, the colorcurve again becomes blue and continues upto 150 d. This marks also the phase where we see a change in the mass-loss rates of the evolution (see subsection 6).
@@
We compare the absolute magnitude (-band) lightcurve of SN 2021foa with a group of SNe IIn and SNe Ibn. For the cases where -band in not available, we use Johnson Cousin -band. The absolute magnitude lightcurve of SN 2021foa behaves similarly with other events having precursor activities. The precursor lightcurve (Event A) had an absolute magnitude 14 mag (R22) similar to SNe 2009ip, 2016jbu and 2021qqp (Hiramatsu et al., 2024). The second peak in the lightcurve (Event B) lies fairly intermediate (MV = 17.8) among the SNe IIn and SNe Ibn (c.f.r Figure 1 of R22). Kiewe et al. (2012), through their studies have found have found that most SNe IIn with precursor events typically rises to second peak around 17 days. The event B typically rises to a maximum with absolute magnitude 18 mag 0.5 (Kiewe et al., 2012; Nyholm et al., 2020) followed by a bumpy decline. Our SN 2021foa also reaches a peak mag at around 20 days, however, SN 2021foa has a low luminosity compared to the sample of Kiewe et al. (2012); Nyholm et al. (2020). Overall, the light curve of SN 2021foa is similar in luminosity to both SNe IIn and SNe Ibn population with L 1042 - 1043 erg sec-1, but the lightcurve resembles more those of SNe IIn given the heterogeneity. The SNe Ibn in our comparison sample have instead more short-lived and less bumpy lightcurves, in accordance with the sample presented by Hosseinzadeh et al. (2017). Also, interestingly, some events with precursor activity like SNe 2009ip and 2018cnf show a light curve shoulder similar to that of SN 2021foa, 20 d post -band maximum. This is most likely associated with the change in the mass-loss rate happening years before explosion. This is also affected by the opacity effects influencing the lightcurve behaviour.